OpenStudy (anonymous):

which decimal is equivalent to 7 and 12/16 A.7.12 B.7.16 C.7.25 D7.75

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

7 12/16=7.75 i can't really explain how to do it

OpenStudy (alexandervonhumboldt2):

jut 12/16=0.75 0.75+7=7.75

OpenStudy (isaiahjwolfe):

You convert a fraction into a decimal by dividing the numerator by the denominator. Then add your decimal (.75) to 7 and you get your answer, D. 7.75